OpenStudy (henriquerb):

Integrate dx/x(1+x^m) with m Real and m differente from 0. The exercise tells to add and subtract x^m on the numerator. I'd love some help. Thank you =)

OpenStudy (anonymous):

\[\frac{1}{x(1+x^m)}=\frac{1+x^m-x^m}{x(1+x^m)}=\frac{1}{x}-\frac{x^{m-1}}{1+x^m}\] Now setting \(u=1+x^m\) gives \(du=mx^{m-1}\,dx\) (for the second term).

OpenStudy (henriquerb):

Thank you very much! =)